OpenStudy (anonymous):

differentiate tan^-1((a+btanx)/(b-atanx))

hartnn (hartnn):

try and put b/a = tan (alpha) then (a+btanx)/(b-atanx) will be of the form tan (...+...)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

let me try by using ur method

OpenStudy (anonymous):

@hartnn -:/

OpenStudy (souvik):

tan^-1{(a/b+tanx)/(1-a/btanx)} lets take \[a/b=\tan \alpha \] tan^-1[{tan(alpha)+tanx}/1-tanx*tan(alpha)}] =tan^-1[tan(x+alpha) =x+alpha

hartnn (hartnn):

oh, it was a/b = tan alpha , not b/a :P

OpenStudy (anonymous):

oh :OOO is there any other method to solve it ???????

OpenStudy (anonymous):

i thought u can only substitute x variable only

hartnn (hartnn):

i feel this is the easiest :) here we replaced one constant, (a/b) with other ...

OpenStudy (anonymous):

oh :O

OpenStudy (anonymous):

@souvik @hartnn -why allpha not tanx ???????

OpenStudy (anonymous):

alpha*

hartnn (hartnn):

tan x ? ''x'' is a variable, whereas a and b are constants we can replace a constant by a constant, and variable by a variable not a constant by a variable and....

OpenStudy (anonymous):

oh :OOO